# Exact Quantum Capacity of Decohering Channels in Arbitrary Dimensions

### v1

@@ 6 Discussion
We have derived the exact quantum capacity for a general class of decohering channels $\Lambda_{k}(\rho)=(1-x)\rho+xD_{k}(\rho)$. While exact quantum capacities are known for certain qubit Pauli channels (e.g., bit-flip with two Kraus operators), their higher-dimensional generalizations typically lack closed-form solutions due to superadditivity. In contrast, our decohering channels $\Lambda_{k}$, though structurally distinct from Pauli models. admit exact closed-form expressions for capacities for all values of $k$, $d$ and $x$.
Our results show that that even under maximal decoherence, ($x=1$) the channel retains non-zero capacity when coherence is preserved within subspaces. The capacity of course depends on the dimension of these subspaces. This challenges the intuition that complete dephasing destroys all quantum information, showing instead that structured noise can leave "islands" of coherence intact. Therefore, survival of quantum capacity under maximal decoherence (when $x=1$), suggests that one can protect information by encoding it into appropriate decoherence-free [24] subspaces.
The capacity formula explicitly quantifies how increasing the subspace dimension $k$ interpolates between classical and quantum transmission limits, with $k=1$ (full dephasing) and $k=d$ (th identity channel) as extremal cases. This dimensional hierarchy provides a practical toolkit for optimizing fault-tolerant encodings: i.e. when the type of coherence is known, one can use suitable encoding to protect quantum information. Our work thus establishes rigorous benchmarks for quantum error correction and underscores the role of subspace-localized coherence as a resource in noisy quantum systems.
Our work opens up several new directions in the problem of capacity of quantum channels. The first one is to use these results for obtaining upper or lower bounds for more general models of noise in arbitrary dimensions, i.e. when incomplete damping occurs in off-diagonal terms. The second avenue is based on the observation that the two crucial properties of this channel, namely $\Lambda\circ D=D$ and $\Lambda^{c}\circ D=\Lambda^{c}$, which have led to degradability of the channel $\Lambda$. One can ask whether there is a systematic way to find other maps ${\cal E}$, and ${\cal N}$, such that ${\cal E}\circ{\cal N}={\cal N}$ and ${\cal E}^{c}\circ{\cal D}={\cal E}^{c}$. One can see if these channels can activate superactivation [41, 42, 43] with zero-capacity channels. Finally The exact solutions we provide open avenues for designing noise-resilient quantum memories and networks, bridging fundamental limits with engineering constraints in the NISQ era [44, 45, 46, 47].
Acknowledgement: V.K. Thanks Dariusz Chruscinski, Saverio Pascazio, Gniewko Sarbicki and Mario Ziman for their valuable comments, when this work was presented in the 56th Symposium in Mathematical Physics in Torun, Poland.

### v2

@@ 6 Discussion
In this work, we have derived an exact, closed-form expression for the quantum capacity of a wide class of decohering channels in arbitrary dimensions. These channels, which interpolate between ideal transmission and complete decoherence, are shown to be degradable for all noise parameters—allowing their capacity to be computed without regularization. This result significantly expands the set of known channels with exactly computable capacities, particularly in higher dimensions where such cases are extremely rare.
Beyond the exact formulas, our findings reveal a deeper structural insight: the geometry of decoherence—specifically the structure of coherence-preserving subspaces—plays a decisive role in determining whether a channel can sustain quantum information transmission, see also [59]. We show that even fully decohering channels may retain nonzero quantum capacity, provided that the decoherence acts in a way that respects overlap between certain subspaces. This challenges the conventional assumption that maximal decoherence necessarily eliminates all capacity.
The channels we analyze may serve as analytically tractable models for noise processes in a variety of physical platforms, including quantum memories, spin chains, and systems with partial access to coherence-preserving degrees of freedom. Our results therefore provide a bridge between the theoretical structure of quantum capacity and experimentally relevant noise models, with implications for designing robust encoding strategies in the presence of structured decoherence. These exact solutions open avenues for designing noise-resilient quantum memories and networks, bridging fundamental limits with engineering constraints in the NISQ era [60, 61, 62, 63]. Finally, our work opens up several new directions in the problem of capacity of quantum channels. The first one is to use these results for obtaining upper or lower bounds for more general models of noise in arbitrary dimensions. The second avenue is based on extending our method for proving degradability to more general channels.
Acknowledgement: We thank Stefano Chessa for his careful reading of the manuscript and for helpful correspondence, including drawing our attention to some relevant works. V.K. Thanks Dariusz Chruscinski, Saverio Pascazio, Gniewko Sarbicki and Mario Ziman for their valuable comments, when this work was presented in the 56th Symposium in Mathematical Physics in Torun, Poland.